OpenStudy (anonymous):

Given u = log(x) and v = log(y). Express log (y^2/^3√x^4) in terms of u and v.

OpenStudy (shiraz14):

Could you please re-write the equation using the Equation function (I presume there is something amiss with the denominator)? Thanks.

OpenStudy (anonymous):

\[\frac{ Y^{2} }{ \sqrt[3]{x^4} }\]

OpenStudy (anonymous):

@shiraz14

OpenStudy (shiraz14):

\[\log(\frac{ y^{2} }{ \sqrt[3]{x^{4}}}) = 2\log(y) - \frac{ 4 }{ 3 }\log(x) = 2v - \frac{ 4 }{ 3 } u\]

OpenStudy (anonymous):

so what exactly did you do?

OpenStudy (shiraz14):

Basically we use the following identities: 1. log(a/b) = log(a) - log(b) 2. log(a^b) = blog(a) And you'd need to convert the root to an exponential - i.e. the cuberoot of x = x^(1/3), for instance.

OpenStudy (anonymous):

alright thanks its it a bit confusing but ill figure it out with some practice

OpenStudy (shiraz14):

YW